Calculate the molar mass of the following substances.
Phosphorus molecule, P4 (Atomic mass of phosphorus = 31)

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

Phosphorus molecule, P4 = 4 × 31 = 124 g

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Molar mass of Phosphorus molecule P4  = 4 × Mass of P
= 4 × 31 = 124 g/mol
 
Hope it helps!
Thank you ✌️